Find the area of the trapezoid
bases are 8 & 4 inches. Non Parallel sides are 5 and 7 inches and height is 7 inches.
Perimeter is 24

Answer by checkley75(3666)

BREAKING THIS TRAP INTO 2 FIGURES OF RECTANGLE & A TRIANGLE WE HAVE:
THERECTANGLE FROM THE MIDDLE IS:
HEIGHT OF 7 & LENGTH OF 4 OR 7*4=28 SQ. INCHES.
NOTE THAT ONE OF THE NON PARALLEL SIDES IS 7 INCHES THE SAME AS THE HEIGHT.
THAT LEAVES THE REMAINING 4 INCHES OF THE LONGER BASE AS THE BASE FOR THE REMAINING TRIANGLE.
WHAT IS LEFT OF ONE END IS A RIGHT TRIANGLE OF HEIGHT = 7 BASE =4 & SIDE OF 5:
THIS AREA=4/2*7=2*7=14 SQ. INCHES.
NOW WE ADD THESE 2 AREAS & WE GET 28+14=42 SQ INCHES.